OpenStudy (anonymous):

if x^2 -3x +4 =0 show that x^4 = 3x-20

OpenStudy (vishweshshrimali5):

\[\large{x^2 - 3x + 4 = 0}\] \[\large{\implies x^2 = 3x-4}\tag{1}\] Now squaring both sides: \[\large{(x^2)^2 = (3x-4)^2}\] \[\large{\implies x^4 = 9x^2 + 16 - 24 x}\tag{2}\] Now again we use the value of \(\large{x^2}\) from equation 1 in equation 2: \[\large{x^4 = 9(3x-4) + 16 - 24x}\] \[\large{\implies x^4 = 27x - 36 + 16 - 24x}\] \[\large{\implies x^4 = 3x - 20}\tag{3}\] If you have a doubt in any step feel free to ask :)

OpenStudy (anonymous):

wow, thanks so much, and can i ask you another question

OpenStudy (vishweshshrimali5):

Sure :)

OpenStudy (anonymous):

Hence, or otherwise, find the equation with roots (a)^4 and (b)^4 if the roots of x^2-3x+4=0 and are a, and b

OpenStudy (vishweshshrimali5):

See an equation with roots \(\alpha\) and \(\beta\) can be written as: \[\large{x^2 - (\alpha+\beta) + \alpha \beta = 0}\]

OpenStudy (anonymous):

yep

OpenStudy (vishweshshrimali5):

If an equation of the format: \[\large{ax^2+bx+c = 0}\] has roots \(x_1\) and \(x_2\) then: \[\large{x_1 x_2 = \cfrac{c}{a}}\] And \[\large{x_1+x_2 = \cfrac{-b}{a}}\]

OpenStudy (anonymous):

yes i understand this

OpenStudy (vishweshshrimali5):

Now our equation is \(x^2-3x+4 = 0\). Thus, \[\large{\alpha \beta = \cfrac{c}{a} = \cfrac{4}{1} = 4}\tag{1}\] And \[\large{\alpha + \beta = \cfrac{-b}{a} = \cfrac{-(-3)}{1} = 3}\tag{2}\]

OpenStudy (anonymous):

yep

OpenStudy (vishweshshrimali5):

Now the equation with roots \(\large{\alpha^4 , ~\beta^4}\) will be: \[\large{x^2 - (\alpha^4 + \beta^4) + (\alpha^4\beta^4) = 0}\] \[\large{\implies x^2 - (\alpha^4 + \beta^4) + (\alpha \beta)^4 = 0}\tag{3}\]

OpenStudy (anonymous):

why can you replace a, as a^4, and b as b^4

OpenStudy (vishweshshrimali5):

This is because you are asking for an equation with roots a^4 and b^4

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 See an equation with roots \(\alpha\) and \(\beta\) can be written as: \[\large{x^2 - (\alpha+\beta) + \alpha \beta = 0}\] \(\color{blue}{\text{End of Quote}}\) When roots were a and b the equation was ^^^ When the roots are a^4 and b^4 then equation will be: \(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 Now the equation with roots \(\large{\alpha^4 , ~\beta^4}\) will be: \[\large{x^2 - (\alpha^4 + \beta^4) + (\alpha^4\beta^4) = 0}\] \[\large{\implies x^2 - (\alpha^4 + \beta^4) + (\alpha \beta)^4 = 0}\tag{3}\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

oh okay then

OpenStudy (anonymous):

so then how do you find a^4 + b^4

OpenStudy (vishweshshrimali5):

Okay then we have: \[\large{\alpha^4 +\beta^4 = (\alpha^2 + \beta^2)^2 - 2(\alpha \beta)^2}\] Also: \[\large{\alpha^2 + \beta^2 = (\alpha+\beta)^2 - 2 \alpha \beta}\]

OpenStudy (vishweshshrimali5):

First you calculate a^2 +b^2 and then calculate a^4 + b^4

OpenStudy (vishweshshrimali5):

\(\color{blue}{\text{Originally Posted by}}\) @vishweshshrimali5 \[\large{\alpha \beta = \cfrac{c}{a} = \cfrac{4}{1} = 4}\tag{1}\] \[\large{\alpha + \beta = \cfrac{-b}{a} = \cfrac{-(-3)}{1} = 3}\tag{2}\] \[\large{\alpha^4 +\beta^4 = (\alpha^2 + \beta^2)^2 - 2(\alpha \beta)^2}\] Also: \[\large{\alpha^2 + \beta^2 = (\alpha+\beta)^2 - 2 \alpha \beta}\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (vishweshshrimali5):

Can you find out the answer now ?

OpenStudy (anonymous):

okay, that was really helpful and thanks again for your help. i got the answer so yeah. :) thanks it means a lot

OpenStudy (vishweshshrimali5):

Your welcome